OpenStudy (anonymous):

Evaluate the logarithmic expression. Do not round the expression until the final answer. When necessary, round to the nearest hundredth. log 0.1 8

OpenStudy (whpalmer4):

Is that supposed to be \[log_{0.1} 8\] If so, use the change of base formula to find the result. \[\log_a b = \frac{\log b}{\log a}\]

OpenStudy (anonymous):

yes an ok